OpenStudy (anonymous):

Solve using substitution: 3y − x − 1 = 0 - 4x = -6y

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Pick one equation to solve for one variable. I suggest the first equation and solve for x because that choice will avoid fractions.

OpenStudy (anonymous):

i tried that but i didnt get the answer i needed

OpenStudy (anonymous):

What equation did you get?

OpenStudy (anonymous):

i got 3y-x=2 idk i cant seem to do because i dont know how

OpenStudy (anonymous):

Go back to the first equation: \[3y-x-1=0\] and solve for x using inverse operations.

OpenStudy (anonymous):

Solve for x means to isolate x.

OpenStudy (anonymous):

Inverse operations means if a number is added, then subtract it from both sides. If the number is being multiplied then divide it from both sides.

OpenStudy (anonymous):

so how do i get rid of y

OpenStudy (anonymous):

You don't get rid of it - you move it to the other side. The idea in this first step is to isolate x.

OpenStudy (radar):

One way would be to add x to both sides of 3y-x-1=0

OpenStudy (anonymous):

so it would be 3y-1=0+x

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

You can drop the 0.

OpenStudy (anonymous):

ok so whats next

OpenStudy (anonymous):

Now, replace the x in the other equation with everything that is x equals.

OpenStudy (radar):